OpenStudy (anonymous):

What do you notice about the vertex and the line of symmetry for this equations y = x2 + 4x + 1? What do you notice about the vertex and the line of symmetry for this equations y = x2 + 4x + 1? @Mathematics

hero (hero):

same x value

OpenStudy (anonymous):

\[\large vertex = (\frac{-b}{2a}, f(\frac{-b}{2a}))\] \[x = \frac{-b}{2a}\] or the x value of the vertex.